OpenStudy (anonymous):

Find the circumference of a circle with a diameter 14 inches

OpenStudy (anonymous):

Answer: 43.96 The formula is: C - = pi d and you just plug in d=14 and pi = 3.14. I'm sure you can solve it from there!

OpenStudy (anonymous):

well i got 106.81 inches?

myininaya (myininaya):

radius=diameter/2 and and Circumferce=2*pi*r=2*pi*(14/2)=14*pi which is approximately what melli got

OpenStudy (anonymous):

I don't know how you got 106.81 :/

OpenStudy (anonymous):

i ment centimeters

OpenStudy (anonymous):

106.76

OpenStudy (anonymous):

s*3.14*17=106.76

OpenStudy (anonymous):

2*

OpenStudy (anonymous):

ops i total did the wrong one